. Demand Curves. ISHO-garment is contemplating a T-shirt advertising promotion. Monthly sales data from T-shirt shops marketing indicate that

            𝑄=1,500–200𝑃


1
Expert's answer
2021-05-31T15:09:45-0400

Give equation for demand, we can get the quantity of T-shirts sold at a given price.

Assume, "P=\$4.5 per \space (T-shirt)"

Quantity sold, "Q=1500-200\times4.5=600(T-shirts)"

Also, given the quantity, price per T-shirt can be calculated.

Assume, "Q=900(T-shirts)"

"900=1500-200P"

"200P=600"

"P=\$3"